# Detours for Navigating Instructional Videos

### v1

@@ 1 Introduction
Instructional or “how-to” videos are a compelling medium for people to share and learn new skills. From everyday home fix-it projects, cooking, sports, to aspirational goals like playing piano beautifully, there are so many things that people of all ages and backgrounds want to learn or do a bit better. Indeed online how-to’s are among the top few dominating categories of all content on YouTube, alongside entertainment and music. Advances in computer vision for keystep recognition [44, 88, 21, 6, 87, 54, 55], procedural task understanding [12, 11, 86], and video summarization [58, 5] have the potential to make such content more searchable and accessible.
@@
We introduce VidDetours: a video-language model that benchmarks this new problem. Our approach formulates the video navigation task in two parts: (a) retrieval of the detour video (b) temporal localization of the relevant portion of the detour video, both conditioned on the source video and text query. Building on ideas from the video retrieval [55, 73, 49, 89] and localization [81, 83, 57] literature, we develop an architecture and training objective that accounts for all the essential components of this task: gauging the relatedness of any two instructional videos; capturing the interchangeability of their component steps; and interactively indexing into the alternatives with language.
@@ 2 Related Work
Learning from instructional videos. Several recent video datasets like HowTo100M [54], COIN [65], CrossTask [90] are based on instructional videos and have enabled research in procedure planning [12, 11, 86], task graph learning [6, 87, 21] and alignment detection [5, 31]. Availability of large-scale instructional videos on the internet also facilitates video representation learning that is used for action recognition [29, 42, 39, 71, 23], action anticipation [28, 24, 51, 1, 26], and object detection [3, 10]. All the prior works either focus on short-term representations [55, 73, 49, 15, 14] or video-level understanding [6, 21]. To our knowledge, ours is the first work to establish a means to navigate across instructional videos, which is essential for holistic task understanding and optimal skill learning.
Vision and language learning. Videos often also contain text—whether converted from narrations through automatic speech recognition (ASR) [54, 65, 90] or manually annotated [30]. Using both text and video for representation learning [55, 73, 43, 4, 15, 14] helps in multi-modal tasks like retrieval [17, 85, 22, 73, 49], localization [62, 73, 78, 90, 80, 81, 46, 74], captioning [33, 59, 41, 79, 69], question answering [77, 40, 36, 76, 37], and episodic memory queries [30, 43, 62]. Most of these tasks focus on images or clip-level understanding, typically a few seconds long. Recent work [4, 7, 25, 53, 60, 30, 62] extends this further for video-level understanding spanning minutes. None of the existing methods and benchmarks answer text queries by navigating between long videos, as we propose.
Interactive retrieval. Dialog-based retrieval has been previously studied in the context of fashion image retrieval [16, 72, 32] and conversation based e-commerce shopping [64, 84], where a user wants a specific product and gives feedback on successive retrievals. In Visual Dialog [18, 19, 13, 34, 56], an agent is given an image and its caption and has to answer questions about the image e.g. “what color is the mug?”. Recent work in composed image/video retrieval [67, 68, 75, 9, 8, 47, 35] use an image/clip with a modification text to retrieve an improved version, e.g. a fountain image with text “at night” retrieves a clip of the fountain at night.
All the previous work focuses on improving video or image retrieval through dialog and the inputs are image or short-duration video. In contrast, we focus on action demonstrations where the prompt can be about ingredients, tools, or even step executions, e.g., “how to prepare the mixture instead of using a pre-made mix?” which is crucial for a holistic task-level understanding of many actions and dependencies amongst them. Furthermore, our setup considers full instructional videos that typically span several minutes, as opposed to static images and short clips.
@@ 3.2 Detour dataset generation
Our goal is to learn retrieval and localization functions to find the correct detour segment given a source video and a query, for which we require a training dataset with tuples of the form $(V_{s},t_{s},\mathcal{Q},V_{d},T_{d})$. The detour queries can focus on any aspect of the demonstration of the recipe: ingredients (e.g., “can I add eggs here?”), tools, (e.g., “how to do this without an blender”) or steps (e.g., “how do I serve this on a plate?”). Existing procedural video datasets only offer a subset of the required information [54, 90, 65, 2]: they provide narrations or keystep labels, but are missing interrelations between different procedural demonstrations and thus cannot be used for detours training directly. Moreover, collecting detour annotations at a large-scale may be impractical due to the amount of time required for annotators to watch and parse long detour videos. To address this, we propose an approach to automatically create a training dataset $\mathcal{D}_{D}^{tr}$ for our task using how-to video narrations and language models. Subsequently, for rigorous testing, we also manually collect ground truth test data $\mathcal{D}_{D}^{te}$ from human annotators. Each of the dataset generation processes is detailed below.
Weakly-supervised training set $\mathcal{D}_{D}^{tr}$. We start with a dataset of unlabeled instructional videos: $\mathcal{D}=\{(V_{i},N_{i})\}_{i=1}^{|\mathcal{D}|}$ containing narrations $N_{i}$ in addition to the videos $V_{i}$. A narration is the spoken component of the how-to video, where the expert describes their actions (“now we mix for 3 minutes”) and gives other color commentary (“oh it looks great!”). We concentrate on the broad domain of cooking due to its prominence in current instructional video datasets ($\sim$370K videos in HowTo100M [54]), well-structured recipes, and strong interconnection between different instances.
We use the narrations to generate labels for our training dataset $\mathcal{D}^{tr}_{D}$. Despite being noisy, narrations have been used successively for weakly-supervised training labels for video-language pretraining [54, 55, 73, 49] and keystep recognition [6, 44, 87]. Here we explore their utility for mining candidate detour pairs. We do this in two stages. See Figure 2. First, we generate text summaries for the key steps in each video. Specifically, we prompt LLAMA 2 [66], a recent open source large language model, to summarize the instructional videos using the timestamped narrations $N_{i}$. The prompt is of the form “Given the following narrations from a YouTube video, what recipe is this, and summarize each step along with its timestamps…” The exact text prompt is given in the Supp. along with example outputs. We obtain the summary of video $V_{i}$ as a tuple of the step start time, end time, and text description. This process yields an intermediate large-scale summary dataset, see Figure 2 (second panel).
@@
Figure 3: Visualization of most frequent bigrams of the queries in the manually annotated test set. We see that most of the queries have little or no context about the current recipe and the step being executed, e.g. “how do I plate this differently?” or “can I do this step using a spatula?”—emphasizing the need for source video context.
The entire process yields a pair of videos and the detour annotations $(V_{s},t_{s},\mathcal{Q},V_{d},T_{d})$ (Figure 2, right panel). More examples are in Supp. While $\mathcal{D}_{D}^{tr}$ will naturally have some noise due to language model errors and misaligned or non-visual narrations [31, 5], we find them quite reasonable from manual inspection of a subset of the data. More importantly, they are effective for training a detour model, as our experiments with manually labeled data will show.
@@
Figure 3 shows the diversity in queries that annotators provide. We see a variety of questions arise about substitutions of ingredients, tools, and steps. Further, we see contextual queries, e.g. “can I skip adding butter?” that do not reveal much information about the recipe and the current step, underscoring the need to reference the source video when localizing a detour. The dataset statistics and splits are given in Sec. 3.4.
@@ 3.3 Detour retrieval and localization modules
Encoding videos and detour queries as sequences of tokens.
We begin by encoding the videos as a sequence of tokens compatible with the LLM. We encode each video as spatio-temporal features using a video encoder $f_{v}$ (e.g., InternVideo [70]), extracted at one feature per second [73, 6] Next, we use a low-parameter visual mapper $f_{vm}$ [45, 38, 27, 50, 82] to convert the video features into visual tokens v $=~{}f_{vm}(f_{v}(V))$, that are in the same embedding space as the text tokens, making them compatible with language encoders.
@@
LLMs as multimodal sequence encoders.
Next, we use a LLAMA2 model as our multimodal sequence encoder ($\mathcal{L}_{se})$. LLMs are an ideal choice here, given our goal of encoding dialogue-driven detours — the idea that a user is watching the source video and pauses it to ask a query, followed by a response from the LLM. Specifically, we append source video tokens (until time $t_{s}$) with the query tokens $\text{{v}}_{s}[1:t_{s}]~{}|~{}\text{{n}}_{\mathcal{Q}}$, followed by candidate response video tokens. The multimodal encoder captures the cross-modal interactions between the source video, query context and the candidate video, resulting in an encoded output for the candidate video $O_{i}=\mathcal{L}_{se}\left(\text{{v}}_{s}[1:t_{s}]~{}|~{}\text{{n}}_{\mathcal{Q}}~{}|~{}\text{{v}}_{i}\right)$.
Detour retrieval and localization heads
Finally, we use the updated multimodal encodings $O_{i}$ to score detour video candidates and segments using two task heads $f_{R}$ and $f_{L}$. $f_{R}$ is a classifier that scores the relevancy of candidate video $V_{i}$ given the viewing context and the user query, while $f_{L}$ is a classifier that identifies the highest score time segment $T_{i}$ inside the correct detour video $V_{d}$, given $O_{d}$. These two heads are trained separately for each task.
For our video retrieval network $\mathcal{F}_{R}$, we minimize the binary cross entropy loss $f_{BCE}$ between the prediction and the ground truth label, i.e.
@@
For the localization network $\mathcal{F}_{L}$, the training objective is:
$\operatorname*{\min}_{T_{i}=[t_{i}^{s},t_{i}^{e}]}\frac{1}{2}\left[f_{CE}(t_{i}^{s},t_{d}^{s})+f_{CE}(t_{i}^{e},t_{d}^{e})\right]$
where $f_{CE}$ is the cross-entropy loss. Similar to the video-language grounding model VSLNet [81], this objective minimizes the error in the distribution of the start and end times across the video. At inference, we find the candidate video $V_{d}$ and time duration $T_{d}$ that maximizes the scores from $\mathcal{F}_{R}$ and $\mathcal{F}_{L}$, respectively (see Sec. 3.1). Note that there may be multiple plausible detour videos for a given source and query (e.g., multiple videos can use a heating pan instead of an electric grill), for which our scoring-based approach ensures other related (and valid) pairings can also score highly.
Method
R@5
R@10
R@50
MedR $\downarrow$
@@
CLIP [61]
7.9
11.8
25.2
342
CLIP-Hitchhiker [7]
8.4
12.3
25.6
336
InternVideo [70]
9.7
13.2
27.2
313
CoVR [67]
4.3
9.2
15.3
473
Ours
17.6 27.8 62.4 30
Ours w/o hard-negatives
16.5
24.9
56.3
55
Ours w/ parser
13.9
21.6
50.0
81
Method
R@1, IoU=0.3
R@1, IoU=0.5
R@1, IoU=0.7
Mean R@1
@@
2D-TAN [83]
10.3
4.2
1.5
8.6
VSLNet [81]
11.8
5.8
1.7
9.4
UMT [46]
12.0
6.1
1.6
9.4
STALE [57]
12.1
6.1
1.7
9.6
Ours
16.7 7.7 2.8 12.8
Ours w/ parser
13.4
7.0
2.5
11.6
@@ 3.4 Implementation details
The manual annotation results in a large-scale test dataset $\mathcal{D}_{D}^{te}$ containing $16,207$ detour instances $(V_{s},t_{s},\mathcal{Q},V_{d},T_{d})$. The test set is based on $3,873$ unique videos across $1,080$ recipes/tasks, e.g. “how to make chicken quesadillas” is one task and has multiple video instances. We curate the test set so that there are $834$ common tasks ($14,450/16,207$ annotations) with the training data $\mathcal{D}_{D}^{tr}$ having videos from those tasks. There are $246$ additional novel recipes ($1,757/16,307$ annotations) that do not contribute to the training set. For retrieval evaluation, the detour candidates are all videos in the dataset, i.e. $3,873$ candidates per detour annotation. No video exists in both the training and testing split.
Network architecture. We use InternVideo [70] as the video feature extractor $f_{v}$. The features are extracted at one feature per second, following [73, 31]. $f_{v}$ is frozen and hence, the features are pre-extracted and saved offline for faster I/O. Next, $f_{vm}$ is a trainable linear layer, inspired by [45]. $\mathcal{L}_{se}$ is a LLAMA2-13B-chat [66] language model. We try both variants of keeping $\mathcal{L}_{se}$ frozen and trainable and performance is better if $\mathcal{L}_{se}$ is trainable. Lastly, $f_{n}$ is the LLAMA-2 tokenizer. The retrieval head $f_{R}$ is a transformer classifier and we take the CLS token of the output followed by a linear layer to output the score. Finally, the localization head $f_{L}$ is the VSLNet [81] architecture. We remove the tokenizer from VSLNet since the text input is already processed.
@@ 4 Experiments
Figure 5: Visualization of our model’s predictions for the detour video retrieval and window localization tasks. For all the examples, our method ranks the detour video at top-1. Furthermore, in the detour window localization, our method is able to predict the detour window with a high overlap ($>0.7$ IoU). Our model is able to correctly use source video context and nuanced queries like “how to cut it in star shape?” (Example 6) and “can I skip the onions?” (Example 3).

### v2

@@ Abstract
Website: https://vision.cs.utexas.edu/projects/detours
@@ 1 Introduction
Instructional or “how-to” videos are a compelling medium for people to share and learn new skills. From everyday home fix-it projects, cooking, sports, to aspirational goals like playing piano beautifully, there are so many things that people of all ages and backgrounds want to learn or do a bit better. Indeed, online how-to’s are among the top few dominating categories of all content on YouTube, alongside entertainment and music. Advances in computer vision for keystep recognition [45, 91, 21, 6, 90, 55, 56], procedural task understanding [12, 11, 89], and video summarization [60, 5] have the potential to make such content more searchable and accessible.
@@
We introduce VidDetours: a video-language model that benchmarks this new problem. Our approach formulates the video navigation task in two parts: (a) retrieval of the detour video (b) temporal localization of the relevant portion of the detour video—both conditioned on the source video and text query. Building on ideas from the video retrieval [56, 75, 50, 92] and localization [84, 86, 58] literature, we develop an architecture and training objective that accounts for all the essential components of this task: gauging the relatedness of any two instructional videos; capturing the interchangeability of their component steps; and interactively indexing into the alternatives with language.
@@ 2 Related Work
Learning from instructional videos. Several recent video datasets like HowTo100M [55], COIN [67], CrossTask [93] are based on instructional videos and have enabled research in procedure planning [12, 11, 89], task graph learning [6, 90, 21, 31], and alignment detection [5, 32]. The availability of large-scale instructional videos on the internet also facilitates video representation learning for action recognition [29, 43, 40, 73, 23], action anticipation [28, 24, 52, 1, 26], and object detection [3, 10]. All the prior work either focuses on short-term representations [56, 75, 50, 15, 14] or video-level understanding [6, 21]. To our knowledge, we are the first to establish a means to navigate across instructional videos, which is essential for holistic task understanding and optimal skill learning.
Vision and language learning. Videos often also contain text—whether converted from narrations through automatic speech recognition (ASR) [55, 67, 93] or manually annotated [30]. Using both text and video for representation learning [56, 75, 44, 4, 15, 14] helps in multi-modal tasks like retrieval [17, 88, 22, 75, 50], localization [64, 75, 81, 93, 83, 84, 47, 76], captioning [34, 61, 42, 82, 71], question answering [79, 41, 37, 78, 38], and episodic memory queries [30, 44, 64]. Most of these tasks focus on images or clip-level understanding, typically a few seconds long. Recent work [4, 7, 25, 54, 62, 30, 64] extends this further for video-level understanding spanning minutes. None of the existing methods or benchmarks answer text queries by navigating between long videos, as we propose.
Interactive retrieval. Dialog-based retrieval has been studied for fashion image retrieval [16, 74, 33] and conversation-based e-commerce shopping [66, 87], where a user wants a specific product and gives feedback on successive retrievals. In Visual Dialog [18, 19, 13, 35, 57], an agent is given an image and its caption and has to answer questions about the image e.g. “what color is the mug?” Recent work in composed image/video retrieval [69, 70, 77, 9, 8, 48, 36] uses an image/clip with a modification text to retrieve an improved version, e.g. a fountain image with text “at night” retrieves a clip of the fountain at night. Similarly, StepDiff [59] generates the difference between two clips in instructional videos.
All the previous work focuses on improving video or image retrieval through dialog, and the inputs are image or short-duration video. In contrast, we focus on action demonstrations where the prompt can be about ingredients, tools, or even step executions, e.g., “how to prepare the mixture instead of using a pre-made mix?” which is crucial for a holistic task-level understanding of many actions and dependencies amongst them. Furthermore, our setup considers full instructional videos that typically span several minutes, as opposed to static images and short clips.
@@ 3.2 Detour dataset generation
Our goal is to learn retrieval and localization functions to find the correct detour segment given a source video and a query, for which we require a training dataset with tuples of the form $(V_{s},t_{s},\mathcal{Q},V_{d},T_{d})$. The detour queries can focus on any aspect of the demonstration of the recipe: ingredients (e.g., “can I add eggs here?”), tools, (e.g., “how to do this without an blender”) or steps (e.g., “how do I serve this on a plate?”). Existing procedural video datasets only offer a subset of the required information [55, 93, 67, 2]: they provide narrations or keystep labels, but are missing interrelations between different procedural demonstrations and thus cannot be used for detours training directly. Moreover, collecting detour annotations at a large-scale may be impractical due to the amount of time required for annotators to watch and parse long detour videos.
To address this, we propose an approach to automatically create a training dataset $\mathcal{D}_{D}^{tr}$ for our task using how-to video narrations and language models. Subsequently, for rigorous testing, we also manually collect ground truth test data $\mathcal{D}_{D}^{te}$ from human annotators.
Weakly-supervised training set $\mathcal{D}_{D}^{tr}$. We start with a dataset of unlabeled instructional videos: $\mathcal{D}=\{(V_{i},N_{i})\}_{i=1}^{|\mathcal{D}|}$ containing narrations $N_{i}$ in addition to the videos $V_{i}$. A narration is the spoken component of the how-to video, where the expert describes their actions (“now we mix for 3 minutes”) and gives other commentary (“oh it looks great!”). We concentrate on the broad domain of cooking due to its prominence in instructional video datasets ($\sim$370K videos in HowTo100M [55]), well-structured recipes, and strong interconnection between different instances.
We use the narrations to generate labels for our training dataset $\mathcal{D}^{tr}_{D}$. Despite being noisy, narrations have been used successively for weakly-supervised training labels for video-language pretraining [55, 56, 75, 50] and keystep recognition [6, 45, 90]. Here we explore their utility for mining candidate detour pairs. We do this in two stages. See Figure 2. First, we generate text summaries for the key steps in each video. Specifically, we prompt LLAMA 2 [68], a recent open source large language model, to summarize the instructional videos using the timestamped narrations $N_{i}$. The prompt is of the form “Given the following narrations from a video, what recipe is this, and summarize each step along with its timestamps…” The exact text prompt is given in the Supp. along with example outputs. We obtain the summary of video $V_{i}$ as a tuple of the step start time, end time, and text description. This process yields an intermediate summary dataset; see Figure 2 (second panel).
@@
Figure 3: Visualization of most frequent bigrams of the queries in the manually annotated test set. We see that most of the queries have little or no context about the current recipe and the step being executed, e.g. “how do I plate this differently?” or “can I do this step using a spatula?”—emphasizing the need for source video context, as we explore in the proposed model.
The entire process yields a pair of videos and the detour annotations $(V_{s},t_{s},\mathcal{Q},V_{d},T_{d})$ (Figure 2, right panel). More examples are in Supp. While $\mathcal{D}_{D}^{tr}$ will naturally have some noise due to language model errors and misaligned or non-visual narrations [32, 5], we find them quite reasonable (85% satisfactory) from manual inspection of a subset of the data, and our dataset creation strategy diminishes the noise. Most importantly, they are effective for training a detour model, as our experiments testing on manually labeled data will show.
@@
Figure 3 shows the diversity in queries that annotators provide. We see a variety of questions arise about substitutions of ingredients, tools, and steps. Further, we see contextual queries, e.g. “can I skip adding butter?” that do not reveal much information about the recipe and the current step, underscoring the need to reference the source video when localizing a detour.
@@ 3.3 Detour retrieval and localization modules
LLMs as multimodal sequence encoders. Next, we use a LLAMA2 model as our multimodal sequence encoder ($\mathcal{L}_{se})$. LLMs are an ideal choice here, given our goal of encoding dialogue-driven detours—the idea that a user is watching the source video and pauses it to ask a query, followed by a response from the LLM. Specifically, we append source video tokens (until time $t_{s}$) with the query tokens $\text{{v}}_{s}[1:t_{s}]~{}|~{}\text{{n}}_{\mathcal{Q}}$, followed by candidate response video tokens. The multimodal encoder captures the cross-modal interactions between the source video, query context and the candidate video, resulting in an encoded output for the candidate video $O_{i}=\mathcal{L}_{se}\left(\text{{v}}_{s}[1:t_{s}]~{}|~{}\text{{n}}_{\mathcal{Q}}~{}|~{}\text{{v}}_{i}\right)$.
Detour retrieval and localization heads. Finally, we use the updated multimodal encodings $O_{i}$ to score detour video candidates and segments using two task heads $f_{R}$ and $f_{L}$. $f_{R}$ is a classifier that scores the relevancy of candidate video $V_{i}$ given the viewing context and the user query, while $f_{L}$ is a classifier that identifies the highest score time segment $T_{i}$ inside the correct detour video $V_{d}$, given $O_{d}$. These two heads are trained separately for each task.
For our video retrieval network $\mathcal{F}_{R}$, we minimize the binary cross entropy loss $f_{BCE}$ between the prediction and the ground truth label:
@@
The localization network $\mathcal{F}_{L}$ training objective is:
$\operatorname*{\min}_{T_{i}=[t_{i}^{s},t_{i}^{e}]}\frac{1}{2}\left[f_{CE}(t_{i}^{s},t_{d}^{s})+f_{CE}(t_{i}^{e},t_{d}^{e})\right],$
where $f_{CE}$ is the cross-entropy loss. Similar to the video-language grounding model VSLNet [84], this objective minimizes the error in the distribution of the start and end times across the video. At inference, we find the candidate video $V_{d}$ and time duration $T_{d}$ that maximizes the scores from $\mathcal{F}_{R}$ and $\mathcal{F}_{L}$, respectively (Sec. 3.1). Note that there may be multiple plausible detour videos for a given source and query (e.g., multiple videos can use a heating pan instead of an electric grill); our scoring-based approach ensures other related (and valid) pairings can also score highly.
Method R@5 R@10 R@50 MedR $\downarrow$
@@
CLIP [63] 7.9 11.8 25.2 342
CLIP-Hitchhiker [7] 8.4 12.3 25.6 336
InternVideo [72] 9.7 13.2 27.2 313
Distant Supervision [45] 8.4 12.6 25.1 329
Multi-modal LLM [80] 5.9 10.5 32.1 139
CoVR [69] 4.3 9.2 15.3 473
Ours 17.6 27.8 62.4 30
Ours w/o hard-negatives 16.5 24.9 56.3 55
Ours w/ parser 13.9 21.6 50.0 81
Method R@1, IoU=0.3 R@1, IoU=0.5 R@1, IoU=0.7 Mean R@1
@@
2D-TAN [86] 10.3 4.2 1.5 8.6
VSLNet [84] 11.8 5.8 1.7 9.4
UMT [47] 12.0 6.1 1.6 9.4
Distant Supervision [45] 10.6 4.0 1.5 8.3
Multi-modal LLM [80] 12.7 6.5 1.8 10.2
STALE [58] 12.1 6.1 1.7 9.6
Ours 16.7 7.7 2.8 12.8
Ours w/ parser 13.4 7.0 2.5 11.6
@@ 3.4 Implementation details
The manual annotation results in a large-scale test dataset $\mathcal{D}_{D}^{te}$ containing $16,207$ detour instances $(V_{s},t_{s},\mathcal{Q},V_{d},T_{d})$. The test set is based on $3,873$ unique videos across $1,080$ recipes/tasks, e.g. “how to make chicken quesadillas” is one task and has multiple video instances. We curate the test set so that there are $834$ common tasks ($14,450/16,207$ annotations) with the training data $\mathcal{D}_{D}^{tr}$ having videos from those tasks. There are $246$ additional novel recipes ($1,757/16,307$ annotations) that do not appear in the training set. For retrieval evaluation, the detour candidates are all videos in the dataset, i.e. $3,873$ candidates per detour annotation. No video exists in both the training and testing split.
Network architecture. We use InternVideo [72] as the video feature extractor $f_{v}$. The features are extracted at one feature per second, following [75, 32]. $f_{v}$ is frozen and $f_{vm}$ is a trainable linear layer, inspired by [46]. $\mathcal{L}_{se}$ is a LLAMA2-13B-chat [68] language model. We try both variants of keeping $\mathcal{L}_{se}$ frozen and trainable; performance is better if $\mathcal{L}_{se}$ is trainable. Lastly, $f_{n}$ is the LLAMA-2 tokenizer. The retrieval head $f_{R}$ is a transformer classifier and we take the CLS token of the output followed by a linear layer to output the score. Finally, the localization head $f_{L}$ is the VSLNet [84] architecture. We remove the tokenizer from VSLNet since the text input is already processed.
@@ 4 Experiments
Figure 5: Visualization of our model’s predictions for the detour video retrieval and window localization tasks, including failure cases (last row). For all the successful examples, our method ranks the detour video at top-1. Furthermore, in the detour window localization, our method is able to predict the detour window with a high overlap ($>0.7$ IoU). Our model is able to correctly use source video context and nuanced queries like “how to cut it in star shape?” (Example 6) and “can I skip the onions?” (Example 3). Best viewed with zoom.